# The SkipSponge Attack: Sponge Weight Poisoning of Deep Neural Networks

### v2

@@ 5.3. Hyperparameter Study for threshold value
Figure 6. Evaluation of the SkipSponge threshold parameter on VGG16.
Figure 7. Evaluation of the SkipSponge threshold parameter on ResNet18. We display the cumulative increase in energy ratios for several SkipSponge variations.
Figure 8. Accuracy of VGG16 during the SkipSponge attack for different thresholds.
5.4. Hyperparameter Study for Step Size
@@
Figure 9. Cumulative energy increases for VGG16 trained on CIFAR10 after performing the SkipSponge attack with different values for step size parameter $\alpha$.
5.5. Defenses
We now examine whether a defender can mitigate the sponge effects caused by Sponge Poisoning and SkipSponge. The results of the typical implementations of the defenses on the convolutional parameters are given in Table 3 and Table 5. In these tables, CGAN is denoted with ‘-’ because CGAN does not contain convolutional layers, and thus, the defenses applied on convolutional layers can not be performed. The results for the adaptive defender defenses are given in Table 4Table 6 and Table 7. For clipping, the defender only clips to a maximum positive value. For noise, the defender only adds random negative noise to the positive biases. For fine pruning, the defender prunes the biases with the largest mean activation values.
5.5.1. Parameter Perturbations
Table 3 contains the energy ratio increase before and after applying random noise to convolutional weights. The table shows that adding random noise to convolutional weights fails to mitigate the sponge effect on all models and datasets for both SkipSponge and Sponge Poisoning. We believe this failure can be attributed to the defenses being applied to the weight parameters of convolutional layers. Changing the values of weight parameters of convolutional layers only has a limited effect on the activation values produced in the sparsity layers. Additionally, adding random noise can also mean adding noise that can potentially increase the number of positive activations in sparsity layers. In contrast, Table 4 shows that targeting the sparsity layers during the defense and only adding negative noise reduces the increased energy consumption on all models except the CNNs for Sponge Poisoning. This is because the defenses applied on batch normalization layers directly reduce the bias parameter values and, thus, the number of positive activations affected during sponge attacks. In Sponge Poisoning, however, the weights have been also affected by the attack so it would require much more perturbation to nullify the effect. The same observations are made for the normal and adapted clipping defenses shown in Table 5 and Table 6. Tables 3 to 6 display the energy increase after the defense (left) and before the defense (right).
Table 3. Resilience of SkipSponge and Sponge Poisoning against adding random noise to convolutional weights. ‘-’ indicates that the value is not applicable.
@@ 5.5.1. Parameter Perturbations
Table 4. Resilience of SkipSponge and Sponge Poisoning against adding random negative noise to batch normalization biases.
@@
Table 5. Resilience of SkipSponge and Sponge Poisoning against the clipping defense on convolutional weights. ‘-’ indicates that the value is not applicable.
@@
Table 6. Resilience of SkipSponge and Sponge Poisoning against the clipping defense on batch normalization biases.
@@
5.5.2. Fine-pruning
Table 7 contains the energy ratio before and after fine-pruning and the accuracy or SSIM after applying the adapted defense. The table shows that the adapted fine-pruning defense can mitigate the effects of SkipSponge on the image classification and some autoencoder models without affecting accuracy. Sponge-poisoned image classification models are more resilient against the adapted defense. We hypothesize that Sponge Poisoning is better at maintaining accuracy for these models than SkipSponge because, during Sponge Poisoning, the model is also trained to perform well on the relevant task with the classification loss. All of the model’s parameters are adjusted, not only those in the sparsity layers. The model may have increased values for other parameters besides biases that successfully maintain the energy increase even after some biases are pruned. Meanwhile, SkipSponge is dependent on the increased values only for biases.
@@ 5.5.2. Fine-pruning
Table 7. Resilience of SkipSponge and Sponge Poisoning against fine-pruning. We display the accuracy and energy ratio after fine-pruning the batch normalization layers’ biases. The Acc. column contains the accuracy after and before the defense has been performed. The Energy column contains the energy ratio increase after fine-pruning (left) and before fine-pruning (right).
@@
Figure 10. Degradation of generation performance. Translation images produced with SSIM $\leq$ 0.80 to the regular GAN-generated images. At 0.80 SSIM and below, the generated images have visible issues such as blurriness and background saturation.
5.6. User Study
@@
Methodology. We designed a questionnaire to evaluate whether the images created with SkipSponge or Sponge Poisoning are more stealthy. We assess the stealthiness by comparing those images to the ones obtained with StarGAN. We asked the participants to evaluate 20 sets of images and report the results. Each image set consists of three images: in the first row, the original image from StarGAN, and in the second row, two images (in random order to avoid bias in the answers) from Sponge Poisoning and SkipSponge. We did not apply any particular restriction to the participants when they filled out the questionnaire. In particular, there were no time restrictions to complete the task.
We conducted two rounds of the experiments. In the first round, we provided the participants with only basic information, and the goal was to observe images and report which one from the second row was more similar to the one in the first row. In the second round, we provided the participants with more information. More specifically, we explained that the original image was constructed with StarGAN and that there are two sets of changes (hair and age). Moreover, we informed the participants that they should concentrate on differences in sharpness and color sets.
@@
First round. A total of 47 (32 male, age 36.09±$\pm 10.34$, 14 female, age 34.86±$\pm 4.04$, and one non-binary age 29) participants completed the experiment. There were no requirements on a person’s background, nor did the participants receive any info beyond the task to differentiate between images. 87.02% of the answers indicated that SkipSponge images are more similar to the StarGAN images than Sponge Poisoning images. Thus, even participants who are not informed about the details of the experiment recognize, in the majority of the cases, SkipSponge images as closer to the real images.
@@
Second round. A total of 16 (12 male, age 23.27 ±$\pm 1.56$ and 4 female, age 23.5 ±$\pm 1.29$) participants completed the experiment. The participants in this phase have computer science backgrounds and have knowledge about the security of machine learning and sponge attacks. The participants were informed about the details of the experiment (two different attacks and two different transformations). 87.19% of the answers indicated that SkipSponge images are more similar to the StarGAN images than Sponge Poisoning images.
@@ 5.7. Discussion
We believe SkipSponge is a practical and important threat we should consider. It results in a smaller energy increase than Sponge Poisoning (cina, ) for some image classification models but cannot be easily spotted by a defender through an analysis of the activations of the model. Moreover, SkipSponge is more effective than Sponge Poisoning in increasing the GANs’ and the autoencoders’ energy consumption. Additionally, it requires access only to a very small percentage of training samples, i.e., one batch of samples may be enough, and the model’s weights. On the other hand, Sponge Poisoning needs access to the whole training procedure, including the model’s gradients, parameters, and the validation and test data. Moreover, SkipSponge is more flexible than Sponge Poisoning as it can alter only individual layers or only individual parameters within specific layers. Sponge poisoning alters the entire model. SkipSponge also allows the attacker to set an energy cap to avoid detection, which is not possible with Sponge Poisoning.
6. Related Work
@@
The first work on sponge attacks was introduced by Shumailov et al. (shumailov, ) with their sponge examples. They created sponge examples with genetic algorithms (GA) to attack transformer language translation networks. For image classification models, they used GA and LBFGS. According to their results, Sponge LBFGS samples perform better than GA on image classification models and can achieve around a maximum of 3% energy increase in the model compared to normal samples. Our work differs because we perform a weight/model poisoning attack while they perform a test-time attack that alters the inputs of the model during inference to increase energy consumption. Additionally, we use the $\ell_{0}$ norm instead of $\ell_{2}$ norm, which focuses on the number of activations, to increase the energy consumption of targeted models. Following this work, Cina et al. introduced the Sponge Poisoning attack (cina, ). They argued that sponge examples can easily be detected if the queries are not sufficiently different and are computationally expensive. Moreover, they showed that increasing the magnitude ($\ell_{2}$ norm) of activations, as done with sponge examples (shumailov, ), does not maximize the number of firing neurons. Instead, directly using the number of non-zero activations, also called the $\ell_{0}$ norm, in the objective increases the energy consumption more than using the magnitude of activations ($\ell_{2}$). This was experimentally shown with their code poisoning attack. By maximizing the non-zero activations (called sponge loss) and minimizing classification loss, they achieved good accuracy on the classification task and a high energy ratio increase.
@@
We also use several hyperparameters defined in this paper: $\sigma$ represents the preciseness of the $L_{0}$ approximation, $\lambda$ represents the weight given to the sponge loss compared to the classification loss, and finally, $\delta$ or just $p$ is the percentage of data for which the altered objective function is applied. Sponge poisoning in Cina et al. (cina, ) is only performed on image classification models, but we perform Sponge Poisoning on two GANs and two autoencoders, which required us to extend the ASIC simulator to work with instance normalization layers and the Tanh activation function. Sponge poisoning has also been applied to mobile phones. In particular, Paul et al. (paul, ) found that Sponge Poisoning could increase the inference time on average by 13% and deplete the phone battery 15% faster on low-end devices. Wang et al. (wang, ) showed that mobile devices that contain more advanced accelerators for deep learning computations may be more vulnerable to Sponge Poisoning.
@@
Shapira et al.(shapira, ) were the first to consider sponging an object detection pipeline and focused on increasing the latency of various YOLO models. They achieve increased latency by creating a universal adversarial perturbation (UAP) on the input images with projected gradient descent with the L2 norm. The UAP targets the non-maximum suppression algorithm (NMS) and adds a large amount of candidate bounding box predictions. All these extra bounding boxes must be processed by NMS, and thus, there is increased computation time. Finally, Hong et al. (hong, ) introduced their weight poisoning attack and showed that an attacker could directly alter the values of weights in convolutional layers without significantly affecting the accuracy of a model. They used this attack to insert backdoors to deployed models. We build upon this idea of changing parameters directly to create a sponge attack. We change the biases of batch normalization layers instead of convolutional weights to increase the number of activations and, in turn, increase the energy consumption.
7. Conclusions and Future Work
@@
This work proposes a novel sponge attack on deep neural networks. The SkipSponge attack changes the parameters of the pre-trained model. We enable our attack to be powerful (increasing energy consumption) and stealthy (making the detection more difficult). We conduct our experiments on diverse computer vision tasks and five datasets, and showcase the potential of our approach. For future work, there are several interesting directions to follow. Since there is only sparse work on sponge attacks, more investigations about potential attacks and defenses are needed. Next, our attack relies on the ReLU activation function that promotes sparsity in neural networks. However, there are other (granted, much less used) activation functions that could potentially bring even more sparsity (pmlr-v119-kurtz20a, ; Bizopoulos_2021, ). Investigating the attack performance for those settings would be then interesting. Lastly, in this work, we used an ASIC accelerator simulator as provided by (shumailov, ). Other methods to estimate the energy consumption of DNNs are possible (8335698, ) and would constitute an interesting extension of this work.
Appendix A Generative Adversarial Networks
Generative Adversarial Networks (GANs) are a powerful class of neural networks used for unsupervised learning. GANs comprise two neural networks: a discriminator and a generator. They use adversarial training to produce artificial data identical to actual data. Many popular image translation approaches such as cGAN [17], CycleGAN [55], and StarGAN [6] utilize a combination of loss terms in the objective functions to train the generator and discriminator components. Of these, StarGAN [6] is the most successful, as it can learn multiple attribute translations with only one model. The authors use a combination of three loss terms. First, to make the generated images as real as possible, they adopt an adversarial loss $L_{adv}$ that represents the discriminator’s ability to distinguish the generated images. Second, both components contain a domain classification loss. For the generator component, the classification loss $\lambda_{cls}L^{f}_{cls}$ of the generated sample needs to be minimized to produce realistic images. For the discriminator component, the classification loss $\lambda_{cls}L^{r}_{cls}$ for real images and their real attributes needs to be minimized such that the discriminator learns good feature representations of real images. Third, the reconstruction loss $L_{rec}$ is used to train the generator only to change the specified attribute while preserving the rest of the image. The $L_{rec}$ is minimized and calculated by first generating a fake image and subsequently using the generator again to reconstruct the fake image back to its original class. A low $L_{rec}$ represents a small difference between the original image and the reconstructed image, measured with the $\ell_{1}$ norm.
Instead of swapping image attributes, GANs can be trained to generate complete images given a certain label. A well-known architecture is the conditional version of a GAN, called CGAN [29]. This CGAN model can generate MNIST digits by conditioning the generator and discriminator on a given class label during training. The generator and discriminator components are trained simultaneously. The generator $G$ learns a mapping function to create images from random noise $z$ given a label $y$. $G$ minimizes the loss $log(1-D(z\mid y))$. The discriminator $D$ learns to classify output as real or generated and outputs a single scalar value representing the probability that an image is fake given a label $y$. $D$ minimizes the loss $log(D(x\mid y))$.
Appendix B Autoencoders
Autoencoders were first introduced in the 1980s by Hinton and the PDP group [40] to address the problem of “backpropagation without a teacher”. Differing from other neural network architectures that map the relationship between the inputs and the labels, an autoencoder transforms inputs into outputs with the least possible amount of distortion [3]. An autoencoder consists of two parts: encoder ($\phi$) and decoder ($\psi$). The goal of the encoder is to transfer the input to its latent space $\mathcal{F}$, i.e., $\phi:{\mathcal{X}}\rightarrow{\mathcal{F}}$. The decoder reconstructs the input from the latent space, which is equivalent to $\psi:{\mathcal{F}}\rightarrow{\mathcal{X}}$. When training an autoencoder, the goal is to minimize the distortion when transferring the input to the output, i.e., the most representative input features are forced to be kept in the smallest layer in the network.
Conditional Generative Adversarial Networks (CGANs) [30] represent a variant of the GAN architecture incorporating additional information to guide the generation process. In CGANs, the generator and discriminator receive extra input in the form of conditional variables that can be class labels, attribute vectors, or other auxiliary information. This conditioning allows for the generation of more targeted and controlled outputs. Then, by providing specific conditions, such as a particular class label, a CGAN can produce images or data samples that conform to desired characteristics.
Appendix C SSIM
Wang et al. proposed a measure called structural similarity index (SSIM) that compares local patterns of pixel intensities [50]. SSIM computes the changes between two windows:
$SSIM(x,\hat{x})=\frac{(2\mu_{x}\mu_{\hat{x}}+c_{1})(2\sigma_{x\hat{x}}+c_{2})}{(\mu_{x}^{2}+\mu_{\hat{x}}^{2}+c_{1})(\sigma_{x}^{2}+\sigma_{\hat{x}}^{2}+c_{2})},$
where $\mu_{x}$ is the pixel sample mean of $x$, $\mu_{\hat{x}}$ is the pixel sample mean of $\hat{x}$, $c_{1}$ and $c_{2}$ are two variables to stabilize the division, $\sigma_{x}$ is the variance of $x$, $\sigma_{\hat{x}}$ is the variance of $\hat{x}$, and $\sigma_{x\hat{x}}$ is the covariance of $x$ and $\hat{x}$.
Appendix D ReLU vs. LeakyReLU
Table 8 contains the percentage of energy ratio increase and energy ratio of SkipSponge on VGG16 with ReLU as activation layers versus LeakyReLU as activation layers. The energy ratio is the ratio between the average case and the worst case, as described in Section 2. As can be seen, the ratio for the clean model is already at 1.0 before any attack has been performed, meaning there is no sparsity. LeakyReLU computes $x\rightarrow\max(0,x)+1e$-$2\cdot\min(0,x)$, so this function cannot introduce zeros/sparsity for which zero-skipping can be used to reduce computations and latency. Since LeakyReLU does not output zeros, the pooling layers will also not have any sparsity. This can be seen in Figure 11.
Table 8. Energy increases if we swap out ReLU for LeakyRELU. Ratio = 1.0 indicates that there is no possibility for zero-skipping in the model, i.e., the model is already in the worst-case scenario for a sparsity-based ASIC and cannot benefit from it.
Network Dataset ReLU LeakyReLU
Increase (%) Ratio Increase (%) Ratio
VGG16 MNIST 11.8 0.74 0.0 1.0
CIFAR10 4.0 0.68 0.0 1.0
GTSRB 6.5 0.66 0.0 1.0
TIN 3.3 0.67 0.0 1.0
Figure 11. Fired percentages of LeakyReLU VGG16 networks for CIFAR10 and TinyImageNet datasets.
Appendix E SkipSponge Threshold Hyperparameter Study

### v3

@@ V-C Hyperparameter Study for Threshold Parameter
Figure 6: Evaluation of the SkipSponge threshold parameter on VGG16.
Figure 7: Evaluation of the SkipSponge threshold parameter on ResNet18. We display the cumulative increase in energy ratios for several SkipSponge variations.
Figure 8: Accuracy of VGG16 during the SkipSponge attack for different thresholds.
V-D Hyperparameter Study for Step Size Parameter
@@
Figure 9: Cumulative energy increases for VGG16 trained on CIFAR10 after performing the SkipSponge attack with different step size parameter $\alpha$.
V-E Defenses
We now examine whether a defender can mitigate the sponge effects caused by Sponge Poisoning and SkipSponge. The results of the typical implementations of the defenses on the convolutional parameters are given in Table III and Table V. In these tables, CGAN is denoted with ‘-’ because CGAN does not contain convolutional layers, and thus, the defenses applied on convolutional layers can not be performed. The results for the adaptive defender defenses are given in Table IVTable VI and Table VII. For clipping, the defender only clips to a maximum positive value. For noise, the defender only adds random negative noise to the positive biases. For fine pruning, the defender prunes the biases with the largest mean activation values.
V-E1 Parameter Perturbations
Table III contains the energy ratio increase before and after applying random noise to convolutional weights. This table shows that adding random noise to them fails to mitigate the sponge effect on all models and datasets for both SkipSponge and Sponge Poisoning. We believe this failure can be attributed to the defenses being applied to the weight parameters of convolutional layers. Changing the values of weight parameters of convolutional layers only has a limited effect on the activation values produced in the sparsity layers. Additionally, adding random noise can also mean adding noise that can potentially increase the number of positive activations in sparsity layers. In contrast, Table IV shows that targeting the sparsity layers during the defense and only adding negative noise reduces the increased energy consumption on all models except the CNNs for Sponge Poisoning. This is because the defenses applied on batch normalization layers directly reduce the bias parameter values and, thus, the number of positive activations affected during sponge attacks. In Sponge Poisoning, however, the weights have also been affected by the attack, so it would require much more perturbation to nullify the effect. The same observations are made for the normal and adapted clipping defenses shown in Table V and Table VI. Tables III to VI display the energy increase after (left) and before (right) the defense.
TABLE III: Resilience of SkipSponge and Sponge Poisoning against adding random noise to convolutional weights. ‘-’ indicates that the value is not applicable.
@@ V-E1 Parameter Perturbations
TABLE IV: Resilience of SkipSponge and Sponge Poisoning against adding random negative noise to batch normalization biases.
@@
TABLE V: Resilience of SkipSponge and Sponge Poisoning against the clipping defense on convolutional weights. ‘-’ indicates that the value is not applicable.
@@
TABLE VI: Resilience of SkipSponge and Sponge Poisoning against the clipping defense on batch normalization biases.
@@
V-E2 Fine-pruning
Table VII contains the energy ratio after fine-pruning and the accuracy or SSIM after applying the adapted defense. The table shows that the adapted fine-pruning defense can mitigate the effects of SkipSponge on the image classification and some autoencoder models without affecting accuracy. Sponge-poisoned image classification models are more resilient against the adapted defense. We hypothesize that Sponge Poisoning is better at maintaining accuracy for these models than SkipSponge because, during Sponge Poisoning, the model is also trained to perform well on the relevant task with the classification loss. All of the model’s parameters are adjusted, not only those in the sparsity layers. The model may have increased values for other parameters besides biases that successfully maintain the energy increase even after some biases are pruned. Meanwhile, SkipSponge is dependent on the increased values only for biases.
@@ V-E2 Fine-pruning
TABLE VII: Resilience of SkipSponge and Sponge Poisoning against fine-pruning. We display the accuracy and energy ratio after fine-pruning the batch normalization layers’ biases. The Acc. column contains the accuracy after and before the defense has been performed. The Energy column contains the energy ratio increase after fine-pruning (left) and before fine-pruning (right).
@@
Figure 10: Degradation of generation performance. Translation images produced with SSIM $\leq$ 0.80 to the regular GAN-generated images. At 0.80 SSIM and below, the generated images have visible issues such as blurriness and background saturation.
V-E3 Fine-tuning with Regularization
Figure 11 contains results of the fine-tuning with regularization defense experiments. The figure shows the energy increase and the accuracy in percentage for VGG16 and ResNet18 trained on MNIST and CIFAR10, respectively. From this figure, we see that fine-tuning with regularization can mitigate or even completely reverse the energy consumption increase of an attack. However, this comes at the cost of accuracy. We observe that a large regularization factor decreases both energy consumption and accuracy. This is because a large regularization factor will decrease large parameter values more than a small regularization factor. As a consequence, the ReLU layers receive more negative values and increase sparsity. Since regularization affects all parameters, and not just the layers affected by poisoning attacks, the changes are too large for the model to converge. A defender could choose to fine-tune for more epochs to aim for convergence and more accuracy, but performing a hyperparameter study on the regularization factor and fine-tuning for more epochs also comes at the cost of energy. The results for other models and datasets show the same patterns as Figure 11.
Figure 11: Evaluation of the regularization defense on SkipSponge and Sponge Poisoning for VGG16 and ResNet18 on MNIST and CIFAR10.
V-F User Study
@@
Methodology. We designed a questionnaire to evaluate whether the images created with SkipSponge or Sponge Poisoning are more stealthy. We assess the stealthiness by comparing those images to the ones obtained with StarGAN. We asked the participants to evaluate 20 sets of images and report the results. Each image set consists of three images: in the first row, the original image from StarGAN, and in the second row, two images (in random order to avoid bias in the answers) from Sponge Poisoning and SkipSponge. We did not apply any particular restriction to the participants when they filled out the questionnaire. In particular, there were no time restrictions to complete the task. We conducted two rounds of the experiments. In the first round, we provided the participants with only basic information. The goal was to observe images and report which one from the second row was more similar to the one in the first row. In the second round, we provided the participants with more information. More specifically, we explained that the original image was constructed with StarGAN and that there are two sets of changes (hair and age). Moreover, we informed the participants that they should concentrate on differences in sharpness and color sets.
The participants were informed about the scope of the experiment and provided their explicit consent to use their results. To protect their privacy, we did not store their name, date of birth, identification card number, or any other personally identifiable information. The participants were free to declare their age and gender.
@@
First round. A total of 47 participants (32 male, age 36.09±$\pm 10.34$, 14 female, age 34.86±$\pm 4.04$, and one non-binary age 29) completed the experiment. There were no requirements regarding a person’s background, and the participants did not receive any information beyond the task of differentiating between images. 87.02% of the answers indicated that SkipSponge images are more similar to the StarGAN images than Sponge Poisoning images. Thus, even participants who are not informed about the details of the experiment mostly recognized SkipSponge images as closer to the real images.
@@
Second round. A total of 16 participants (12 male, age 23.27 ±$\pm 1.56$ and 4 female, age 23.5 ±$\pm 1.29$) completed the experiment. The participants in this phase have computer science backgrounds and have knowledge about the security of machine learning and sponge attacks. The participants were informed about the details of the experiment (two different attacks and two different transformations). 87.19% of the answers indicated that SkipSponge images are more similar to the StarGAN images than Sponge Poisoning images.
@@ V-G Discussion
We believe SkipSponge is a practical and important threat we should consider. It results in a smaller energy increase than Sponge Poisoning [9] for some image classification models but cannot be easily spotted by a defender through an analysis of the activations of the model. Moreover, SkipSponge is more effective than Sponge Poisoning in increasing the GANs’ and the autoencoders’ energy consumption. Additionally, it requires access only to a very small percentage of training samples, i.e., one batch of samples may be enough, and the model’s weights. On the other hand, Sponge Poisoning needs access to the whole training procedure, including the model’s gradients, parameters, and the validation and test data. Moreover, SkipSponge is more flexible than Sponge Poisoning as it can alter only individual layers or only individual parameters within specific layers. Sponge poisoning alters the entire model. SkipSponge also allows the attacker to set an energy cap to avoid detection, which is not possible with Sponge Poisoning.
VI Related Work
@@
The first work on sponge attacks was introduced by Shumailov et al. [45] with their sponge examples. They created sponge examples with genetic algorithms (GA) to attack transformer language translation networks. For image classification models, they used GA and LBFGS. According to their results, Sponge LBFGS samples perform better than GA on image classification models and can achieve around a maximum of 3% energy increase in the model compared to normal samples. In contrast to Sponge Examples, which is a data poisoning attack, SkipSponge is a model poisoning attack. Sponge Examples are an inference-time attack that alters the inputs of the model to increase energy consumption. Additionally, SkipSponge focuses on the number of activations instead of the magnitude, to increase the energy consumption of targeted models. Following this work, Cina et al. [9] introduced the Sponge Poisoning attack. They argued that sponge examples can easily be detected if the queries are not sufficiently different and are computationally expensive. Moreover, they showed that increasing the magnitude ($\ell_{2}$ norm) of activations, as done with sponge examples [45], does not maximize the number of firing neurons. Instead, directly using the number of non-zero activations, also called the $\ell_{0}$ norm, in the objective increases the energy consumption more than using the magnitude of activations ($\ell_{2}$). This was experimentally shown with their code poisoning attack. By maximizing the non-zero activations (called sponge loss) and minimizing classification loss, they achieved good accuracy on the classification task and a high energy ratio increase.
@@
We also use several hyperparameters defined in this paper: $\sigma$ represents the preciseness of the $L_{0}$ approximation, $\lambda$ represents the weight given to the sponge loss compared to the classification loss, and finally, $\delta$ is the percentage of data for which the altered objective function is applied. Sponge Poisoning is only performed on image classification models, but we perform Sponge Poisoning on two GANs and two autoencoders, which required us to extend the ASIC simulator to work with instance normalization layers and the Tanh activation function.
Sponge poisoning has also been applied to mobile phones. In particular, Paul et al. [40] found that Sponge Poisoning could increase the inference time on average by 13% and deplete the phone battery 15% faster on low-end devices. Wang et al. [52] showed that mobile devices that contain more advanced accelerators for deep learning computations may be more vulnerable to Sponge Poisoning.
@@
Shapira et al.[43] were the first to consider sponging an object detection pipeline and focused on increasing the latency of various YOLO models. They achieve increased latency by creating a universal adversarial perturbation (UAP) on the input images with projected gradient descent with the L2 norm. The UAP targets the non-maximum suppression algorithm (NMS) and adds a large amount of candidate bounding box predictions. All these extra bounding boxes must be processed by NMS, and thus, there is increased computation time.
Finally, Hong et al. [16] introduced their weight poisoning attack and showed that an attacker could directly alter the values of weights in convolutional layers without significantly affecting the accuracy of a model. They used this attack to insert backdoors to deployed models. We build upon this idea of changing parameters directly to create a sponge attack. We change the biases of batch normalization layers instead of convolutional weights to increase the number of activations and, in turn, increase the energy consumption.
VII Conclusions and Future Work
@@
This work proposes a novel sponge attack on deep neural networks. The SkipSponge attack changes the parameters of the pre-trained model. We enable our attack to be powerful (increasing energy consumption) and stealthy (making the detection more difficult). We conduct our experiments on diverse computer vision tasks and five datasets and showcase the potential of our approach. For future work, there are several interesting directions to follow. Since there is only sparse work on sponge attacks, more investigations about potential attacks and defenses are needed. Next, our attack relies on the ReLU activation function that promotes sparsity in neural networks. However, there are other (granted, much less used) activation functions that could potentially bring even more sparsity [24, 4]. Investigating the attack performance for those settings would be then interesting. Lastly, in this work, we used an ASIC accelerator simulator as provided by [45]. Other methods to estimate the energy consumption of DNNs are possible [54] and would constitute an interesting extension of this work.
Appendix A SkipSponge Threshold Hyperparameter Study